# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

### v1

@@ 3.6.3 Concentration of GVI Posteriors under Arbitrary Divergences
Make Assumptions 1, 2, 5, 6 and 7. Suppose also that there exists a sequence $(\nu_{n})\subset\mathcal{Q}$, not necessarily GVI minimisers, for which (i) $(J_{L_{n}}(\nu_{n})-L_{n}(\Theta))\rightarrow 0$ $P_{0}$–a.s., (ii) $n^{-1}D(\nu_{n}:\Pi)\rightarrow 0$, and (iii) $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ it almost surely holds that $\liminf_{n}L_{n}(A)>L(\Theta)$. Then, for all such sets $A$,
$Q_{n}A\rightarrow 0$
with $P_{0}$–probability 1.
@@
4 Illustrative Examples
We provide worked examples in support of the theory.
4.1 Existence of Everywhere Bounded Divergences
@@
The prime example of such divergences, that satisfy Assumption 3 for unrestricted choices of $\mathcal{Q}$ and $\mathcal{G}$, is the Total Variation Distance defined as:
$D_{TV}(Q:\Pi):=\sup_{A\in\mathcal{T}}\left|QA-\Pi A\right|\leq 1.$ (6)
@@
As an immediate consequence, the Hellinger Distance is also bounded above since by a well known result, see for instance [29, Lemma B.5],
$D_{H}^{2}(Q:\Pi)=\int_{\Theta}\left(\sqrt{\frac{dQ}{d\mu}(\theta)}-\sqrt{\dfrac{d\Pi}{d\mu}(\theta)}\right)^{2}d\mu(d\theta)\leq D_{TV}(Q:\Pi).$ (7)
Furthermore, such a divergence can be constructed as part of the class of ff–divergences [1, 3]. We follow [3] and define these as
@@
Making use of a result in [10], we know that for ff–divergences, there is an upper bound whenever the following limit exists and is finite:
$0\leq D_{f}(Q:\Pi)\leq\lim_{u\rightarrow 0^{+}}\left\{f(u)+uf\left(\frac{1}{u}\right)\right\}$ (8)
@@
4.2 Gaussian Location Model
Continuing from Section 2.3, we use the example of a Gaussian likelihood, and some Gaussian prior with fixed unit variance but misspecified mean. For simplicity, we assume that $X_{1}^{n}\overset{iid}{\sim}P_{0}:=\mathcal{N}(\theta_{0},1)$ and that the variance is known. In terms of densities, to estimate this normal location model, we place a Gaussian prior $\pi(\theta)\sim\mathcal{N}(\mu_{\pi},1)$ over the possible location parameters in $\Theta=\mathbb{R}$, and we assume that the likelihood hypothesis is $p_{\theta}(X)\sim\mathcal{N}(\theta,1)$. The standard loss for Bayesian inference is then naturally
$L(X_{1}^{n},\,\theta)=-\frac{1}{n}\log p_{\theta}(X_{1}^{n})=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(X_{i})$
which $P_{0}$–a.s. converges in $n$ to its limit, $L(\theta)=\mathbb{E}_{X\sim P_{0}}[-\log p_{\theta}(X)]$. The distribution $P^{\star}_{n}$ in $\mathcal{Q}$, the set of Gaussian measures, is then given by:
$p^{\star}_{n}(\theta)=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\mathbb{E}_{q(\theta)}\left[-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(X_{i})\right]$
which we can evaluate for some $q\sim\mathcal{N}(\mu,\sigma^{2})$ as
$-\frac{1}{n}{{\int}}q(\theta)\left(\sum_{i=1}^{n}\log\left[\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{(X_{i}-\theta)^{2}}{2}\right\}\right]\right)\,d\theta$
$=\underbrace{{{\int}}q(\theta)\frac{1}{2}\log 2\pi\,d\theta}_{=\frac{1}{2}\log 2\pi}+\frac{1}{n}\int q(\theta)\underbrace{\left(\sum_{i=1}^{n}\frac{(X_{i}-\theta)^{2}}{2}\right)}_{=\frac{{\sum_{i}X_{i}^{2}-2\theta\sum_{i}X_{i}+n\theta^{2}}}{2}}\,d\theta$
$=\frac{1}{2}\log 2\pi+\sum_{i=1}^{n}\frac{X_{i}^{2}}{2n}-\frac{\mu\sum_{i=1}^{n}X_{i}}{n}+\frac{(\mu^{2}+\sigma^{2})}{2}$
which is convex in both $\mu$ and $\sigma$, and the minimiser corresponds to the degenerate Gaussian centred at the empirical mean, $\overline{X_{1}^{n}}$. So we let $P^{\star}_{n}(\theta)=\delta_{\overline{X_{1}^{n}}}(\theta)$, and find that
$J_{L_{n}}(P_{n}^{\star})=\frac{1}{2}\log 2\pi+\frac{1}{2n}\sum_{i=1}^{n}X_{i}^{2}-\frac{1}{2}\left(\overline{X_{1}^{n}}\right)^{2}.$
We can then find the set $\mathcal{R}^{\star}_{n}$ in which the GVI posterior may lie.
$\int_{\Theta}L(X_{1}^{n},\,\theta)\left[Q(d\theta)-P_{n}^{\star}(d\theta)\right]\leq\frac{M}{\beta n}$
$\implies\left({\frac{1}{2}\log 2\pi}+\sum_{i=1}^{n}\frac{X_{i}^{2}}{2n}-\frac{\mu\sum_{i=1}^{n}X_{i}}{n}+\frac{(\mu^{2}+\sigma^{2})}{2}\right)-J_{L_{n}}(P_{n}^{\star})\leq\frac{M}{\beta n}$
$\implies\left(\frac{1}{2}\overline{X_{1}^{n}}-\mu\right)\overline{X_{1}^{n}}+\frac{\mu^{2}+\sigma^{2}}{2}\leq\frac{M}{\beta n}$ (9)
We note that the LHS is strictly non-negative, and by elementary analysis this parabola has at least one real root, therefore:
$\mu^{2}-2\mu\overline{X_{1}^{n}}-\frac{2M}{\beta n}+\left(\sigma^{2}+\overline{X_{1}^{n}}^{2}\right)\leq 0$
$\implies$ $\mu\in\overline{\mathrm{B}\left(\overline{X_{1}^{n}},\;\left(\frac{2M}{\beta n}-\sigma^{2}\right)\right)}$ (10)
$\implies$ $\sigma^{2}\in\left[0,\;\mu^{2}-2\mu\overline{X_{1}^{n}}+\overline{X_{1}^{n}}^{2}-\frac{2M}{\beta n}\right]$ (11)
where $\overline{\mathrm{B}(x,r)}$ denoted the closed ball of radius $r$ centred at $x$. Hence, the space of possible GVI posteriors is compact regardless of the prior mean $\mu_{\pi}$, in difference to the Bayesian posterior found in Section 2.3.
Furthermore, since the LHS in Equation 9 is non–negative, it is straightforward to see that we converge to the mean of the DGP as $n\rightarrow\infty$, i.e. $\lim_{n}\frac{1}{n}\sum_{i}^{n}X_{i}=\mathbb{E}_{P_{0}}[X]=\theta_{0}$, with parameters $\mu_{\infty}=\theta_{0}$ and $\sigma_{\infty}^{2}=0$ $P_{0}$ almost surely.
This happens irrespective of the prior, even in the case where $\pi(\theta)=\delta_{\theta_{\pi}}$ and $\theta_{\pi}\rightarrow\infty$, or when $\theta_{0}\notin\operatorname*{supp}(\Pi)$. This result stands in contrast to concentration results of GBI posteriors such as Theorem 2 of [28].
4.2.1 A Comparison with Bayesian Inference
To highlight this relevance, we compare the set $\mathcal{R}^{\star}_{n}$ with the Bayesian posterior under prior misspecification and examine in which set the prior may lie in order to produce the worst case posterior found in $\mathcal{R}_{n}^{\star}$.
From Section 2.3, we know that the Bayesian posterior mean and variance are given as
$\sigma_{B}^{2}=\left(\sigma_{\pi}^{-2}+n\right)^{-1},\,$ $\mu_{B}=\frac{1}{\left(\sigma_{\pi}^{-2}+n\right)}\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}{X_{i}}\right)$
Therefore, by Equation 10, supposing that $M=\beta$, and rearranging we have that the prior mean is required to be in the closed ball around the empirical mean
$\mu_{\pi}\in\overline{\mathrm{B}\left(\overline{X_{1}^{n}},\frac{2}{n}+\sigma_{\pi}^{2}\right)}$
in order for the Bayesian posterior to perform similarly to the GVI posterior in terms of robustness to prior misspecification. Equation 11 would make the computation tedious without improving clarity; for illustrative purposes this suffices. Here, we can interpret this as requiring the prior to have its mean within one variance of the empirical mean plus an error term depending on the number of observations.
5 Proofs
We now provide the proofs of the theorems presented in Section 3.
5.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
We first present the following auxiliary lemmata that will be instrumental in the proof of Theorem 1.
@@ 5.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
Let Assumptions 1, 2 and 3 hold. Suppose $\exists C\in\mathbb{R}$ and that $\exists P,Q\in\mathcal{Q}$ such that
then $\forall\Pi\in\mathcal{G}$ it holds that $T_{n}(P)<T_{n}(Q)$.
Proof of Lemma 1.
Recall that $T_{n}(\nu)=n\mathbb{E}_{\nu}[L(X_{1}^{n},\theta)]+\frac{1}{\beta}D(\nu:\Pi)$ for some $\nu\in\mathcal{Q}$, then for any $P,Q\in\mathcal{Q}$ satisfying the assumption above, we have that:
hence $T_{n}(P)<T_{n}(Q)$ for all priors $\Pi\in\mathcal{G}$. ∎
The next lemma is more general and does not require any assumptions on the divergence.
Lemma 2.
Make Assumptions 1, 2 and 5. Suppose $\mathcal{Q}$ is closed with respect to weak convergence. Pick any $C\in\mathbb{R}$, then the subspace $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq C\}$ is tight.
@@
As the loss is lower bounded and lower semi–continuous it achieves a minimum on any compact set, so $\inf_{\theta\in K}L_{n}(\theta)=c\in\mathbb{R}$ is achieved for some $\theta$ and exists, and hence by the lower bound, there exists some global minimum $c^{\star}:=\inf_{\theta\in\Theta}L_{n}(\theta)$. Now, without loss of generality we may assume $c^{\star}<0$ otherwise we may pick 0 as a lower bound instead of $c^{\star}$ below, so that
$\nu K\inf_{\theta\in K}L_{n}(\theta)+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\geq c^{\star}+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\,.$
@@
$J_{L_{n}}(\nu)\geq c^{\star}+\varepsilon\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)$
and additionally, the set $c^{\star},\frac{C-c^{\star}}{\varepsilon}]\subset\mathbb{R}$ is compact. Therefore, by the coerciveness of the loss there exists $K^{\prime}\subset\Theta$ compact such that $L_{n}:\Theta\backslash K^{\prime}\rightarrow\mathbb{R}\backslash[c^{\star},\frac{C-c^{\star}}{\varepsilon}]$, in words there exists some compact set $K^{\prime}$ such that the loss defined on its complement maps $\theta$ exclusively outside of $c^{\star},\frac{C-c^{\star}}{\varepsilon}]$. Note that $K^{\prime}$ does not depend on $\nu$. We may now define $K=K^{\prime}$ to see that if $\mathcal{A}$ is not tight then $\exists\nu\in\mathcal{A}$ for which $J_{L_{n}}(\nu)>C$. As $\varepsilon$ was arbitrary, the claim follows. ∎
Using these lemmata, we are now able to prove Theorem 1.
@@
By Assumption 4, $\exists P_{n}^{\star}\in\mathcal{Q}$ such that
$P_{n}^{\star}\in\operatorname*{arg\,inf}_{Q\in\mathcal{Q}}\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)].$
Now let $C=\mathbb{E}_{P_{n}^{\star}}[L(X_{1}^{n},\,\theta)]$. Then for all $Q\in\mathcal{Q}$ that satisfy
$\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]>C+M(n\beta)^{-1},$
by Lemma 1 it must hold that $T_{n}(P_{n}^{\star})<T_{n}(Q)$, and hence $Q(d\theta)$ is not a minimiser of the GVI objective. Therefore, the only measures $Q^{\prime}\in\mathcal{Q}$ that are potential minimisers of the GVI objective for any prior $\Pi\in\mathcal{G}$ are those that satisfy:
$T_{n}(Q^{\prime})\leq T_{n}(P_{n}^{\star}).$

### v2

@@ 4.3. On Consistency of GVI Posteriors under Arbitrary Divergences
Make Assumptions 1, 2, 3, 4 and 6. Suppose also that there exists a sequence $(\nu_{n})\subset\mathcal{Q}$, not necessarily GVI minimisers, for which (i) $(J_{L_{n}}(\nu_{n})-L_{n}(\Theta))\rightarrow 0$ $P_{0}$–a.s., (ii) $n^{-1}D(\nu_{n}:\Pi)\rightarrow 0$, and (iii) $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ it almost surely holds that $\liminf_{n}L_{n}(A)>L(\Theta)$. Then, for all such sets $A$, $Q_{n}A\rightarrow 0$ with $P_{0}$–probability 1.
@@
5. Illustrative Example
5.1. Mitigating Inconsistent Bayes Posteriors
Theorems 2 and 4 suggest that consistency holds independent of the choice of prior. Hence, we may mitigate inconsistency of Bayes posteriors through GVI with bounded divergences. To illustrate this, we follow the example of Diaconis and Freedman [diaconis1986b] in modelling a location parameter $\theta_{0}$ given independent errors ${\varepsilon}_{i}$. We refer to their model as DF.
(5) $X_{i}=\theta_{0}+{\varepsilon}_{i}$
for some ${\varepsilon}_{i}$ independent with unknown distribution $F_{0}$. In particular, they suppose that it has some density $h_{0}\in C_{c}^{\infty}(\mathbb{R})$ symmetric about 0 which is a strict global maximum. DF model the prior for $\theta_{0}$ and $F_{0}$ independent, $\theta\sim\mathcal{N}(0,1)$ and $F\sim D(\alpha)$ the Dirichlet process prior based on the standard Cauchy.
Notably, DF prove that there exist $h_{0}$, symmetric about $\theta_{0}$, for which it $P_{0}$–a.s. holds that the Bayes posterior does not concentrate on $\theta_{0}=0$ but rather around $\pm\gamma$ for $\gamma>0$ with equal probability. In particular, DF demonstrate $\forall\eta>0$ it holds that $\limsup\Pi_{n}\{(\theta,F):|\theta-\gamma|<\eta\}=1\,a.e.$ and $\limsup\Pi_{n}\{(\theta,F):|\theta+\gamma|<\eta\}=1\,a.e.$ That is the posterior oscillates around $\gamma$ and $\gamma$. In fact, DF show that this holds even if we forgo identifiability issues associated with $\theta$ and $F$ by considering only symmetric densities $G$ given a symmetrised Dirichlet prior.
For our purposes, we consider the location parameter $\theta\in{\mathbb{R}}$ to have corresponding Dirac measure $\delta_{\theta}$, and consider $G$ to be in the set of symmetric probability measures. That is $G\in\overline{{\mathcal{P}}}:=\{\frac{1}{2}(\nu+\nu^{-}):\nu\in{\mathcal{P}}({\mathbb{R}}),\nu^{-}:=\nu(-dx)\}$, which are mean zero, symmetric measures. We consider the set of models to be their convolution $\delta_{\theta}\star G$, that is $\{\delta_{\theta}\star G:\theta\in{\mathbb{R}},G\in\overline{{\mathcal{P}}}\}$ with $\Theta={\mathbb{R}}\times\overline{{\mathcal{P}}}$.
We now demonstrate in Proposition 8 that GVI updating with bounded divergences results in consistent posteriors.
Proposition 8.
Suppose the model is given as Equation 5 with ${\varepsilon}_{i}$ independent, $h_{0}\in C^{\infty}_{c}$ symmetric about zero and such that the Bayes posterior is inconsistent, and $\theta_{0}=0$ as in [diaconis1986b, Theorem 3]. Let $\pi(d\theta,dG)$ be as above, $L(X_{1}^{n},\,(\theta,G))$ the average negative log likelihood, and $\mathcal{Q}=\mathcal{P}(\Theta)$ for $\Theta=\mathbb{R}\times\overline{{\mathcal{P}}}$. Then, under any divergence satisfying Assumptions 4 and 5, the GVI posteriors will be consistent.
Moreover, it is $P_{0}$–a.s. that $\forall\eta>0$
$\lim_{n\rightarrow\infty}Q_{n}\{\theta:|\theta|<\eta\}=1\,\,a.e.$
We give explicit examples of divergences satisfying both Assumptions 4 and 5 in Section 5.2. Note also that there exists an equivalent $N_{\varepsilon}$ to $N_{\eta}:=\{\theta\in\mathbb{R}:|\theta|<\eta\}$.
Proof.
We verify the assumptions of Theorem 2. By [kechris1995, Thm. 17.23, Prop. 3.3] we know that $\Theta=\mathbb{R}\times\overline{{\mathcal{P}}}$ is Polish. Assumptions 5 and 4 are satisfied by assumption, and that $L(X_{1}^{n},\,\theta)$ are $\sigma(X_{1}^{n})\times\mathcal{T}$ measurable is immediate, hence Assumption 1 holds. Since $\mathcal{Q}=\mathcal{P}(\Theta)$, we trivially include Dirac measures and Assumption 2 holds.
We find it more intuitive to work with a log likelihood ration in the remainder rather than the negative log likelihood. This is equivalent through addition of a constant term depending on $n$ without affecting any minimisers. Let
$L(X_{1}^{n},(\theta,g))=-\frac{1}{n}\log\frac{p(X_{1}^{n}|\theta,g)}{p_{0}(X_{1}^{n})}\rightarrow$ $\mathbb{E}_{P_{0}}\left[-\log\frac{p(X|\theta,g)}{p_{0}(X)}\right]$
$=D_{KL}(P_{0}:\delta_{\theta}\star G)=:L(\theta,g).$
As $p_{0}(X_{1}^{n})\in(0,\infty)$ almost surely and $n^{-1}\log p(X_{1}^{n}|\theta,g)\in[0,\infty]$, the losses are lower bounded (a.s.). These converge pointwise $P_{0}$–a.s. by the Asymptotic Equipartition Property as $P_{0}$ is stationary and ergodic under the usual shift [algoet1988]. In particular, as the model is well specified, we may pick $\theta=\theta_{0}$ and $g\equiv h_{0}$ to achieve $L(X_{1}^{n},(\theta_{0},h_{0}))=0$ for each $n$, and hence
$\limsup_{n}\inf_{(\theta,g)\in\Theta}L(X_{1}^{n},(\theta,g))\leq 0\leq\inf_{(\theta,g)\in\Theta}L(\theta,g).$
We therefore satisfy Assumption 6. Since we consider symmetric densities $g$, $h_{0}$ is continuous, and the space of bounded continuous functions is dense in $L^{1}(m)$, we may restrict ourselves to continuous densities. It follows that the convolution $p(X_{i}|\theta,g)=\delta_{\theta}\star g(X_{i})$ is continuous and coercivity is immediate as $p(X_{1}^{n}|\theta,g)$ vanishes at the boundary. Assumption 3 is thereby satisfied.
It remains to be shown that $\forall A\subset\Theta$ for which $L(A)>L(\Theta)$ it holds that $\liminf L_{n}(A)>L(\Theta)$. As the model is well specified $L(\Theta)=0$ which holds iff $P_{0}=P_{\theta}$ in the sense of measures. Note that the set $N_{\eta}^{c}$ is such a set as $h_{0}(\theta)$ has a strict global maximum at 0. For the conclusion it is sufficient to verify the condition for all $\eta>0$. Note $p(X_{1}^{n}|\theta,g)=\prod_{i=1}^{n}g(X_{i}-\theta)$. Now since the pairs $(\theta,g)$ are symmetric densities centred at $\theta$, and restricting $|\theta|\geq\eta$ it is immediate that the condition holds, as eventually $P_{0}$–a.s. for all $n$ sufficiently large $p_{0}(X_{1}^{n})>\prod_{i}g(X_{i}-\theta)$ for a.e. $\theta\in N_{\eta}$. ∎
5.2. The Choice of Divergence
@@
Many choices of divergence satisfying Assumptions 5 and 4 exist. The prime example of such divergences, that moreover satisfies Assumption 5 for unrestricted choices of $\mathcal{Q}$ and $\mathcal{G}$, is the Total Variation Distance defined as:
(6) $D_{TV}(Q:\Pi):=\sup_{A\in\mathcal{T}}|QA-\Pi A|\leq 1.$
@@
As an immediate consequence, the Hellinger Distance is also bounded above. By a well known result, see for instance [pinski2015, Lemma B.5],
(7) $D_{H}^{2}(Q:\Pi)=\frac{1}{2}\int_{\Theta}\left(\sqrt{\frac{dQ}{d\mu}(\theta)}-\sqrt{\dfrac{d\Pi}{d\mu}(\theta)}\right)^{2}\mu(d\theta)\leq D_{TV}(Q:\Pi).$
Furthermore, such a divergence can be constructed as part of the class of ff–divergences [ali1966, amari2016]. We follow [amari2016] and define these as
@@
Making use of a result in [amari2016], we know that for ff–divergences, there is an upper bound whenever the following limit exists and is finite:
(8) $0\leq D_{f}(Q:\Pi)\leq\lim_{u\rightarrow 0^{+}}\left\{f(u)+uf\left(\frac{1}{u}\right)\right\}$
@@
Furthermore, Theorem 4 for instance allows for the following divergence:
$D_{TV}^{(M\sqrt{n})}(Q:\Pi):=n^{1/2}M||Q-\Pi||_{TV},$
where $M>0$ is some sufficiently large constant, preventing over–concentration around the data in cases where the prior is not severely misspecified while maintaining the results on rates of convergence. We may hence achieve an interplay between Bayesian and frequentist inference.
6. Proofs
6.1. Existence and Uniqueness of a Generalised Variational Inference Posterior
@@
Make Assumptions 1, 2 and 3. Suppose $\mathcal{Q}$ is closed with respect to weak convergence. Pick any $C\in\mathbb{R}$, then the subspace $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq C\}$ is tight.
@@
As the loss is lower bounded and lower semi–continuous it achieves a minimum on any compact set, so $\inf_{\theta\in K}L_{n}(\theta)=c\in\mathbb{R}$ is achieved for some $\theta$ and exists, and hence by the lower bound, there exists some global minimum $c^{\star}:=\inf_{\theta\in\Theta}L_{n}(\theta)$. Now, take $c^{\star}\wedge 0:=\min\{c^{\star},0\}$, so that
$\nu K\inf_{\theta\in K}L_{n}(\theta)+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\geq(c^{\star}\wedge 0)+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\,.$
@@
$J_{L_{n}}(\nu)\geq(c^{\star}\wedge 0)+\varepsilon\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\,.$
And as the set $c^{\star}\wedge 0,\frac{C-(c^{\star}\wedge 0)}{\varepsilon}]\subset\mathbb{R}$ is compact, by the coerciveness of the loss, there exists $K^{\prime}\subset\Theta$ compact such that $L_{n}:\Theta\backslash K^{\prime}\rightarrow\mathbb{R}\backslash[c^{\star}\wedge 0,\frac{C-(c^{\star}\wedge 0)}{\varepsilon}]$, in words there exists some compact set $K^{\prime}$ such that the loss defined on its complement maps $\theta$ exclusively outside of $c^{\star}\wedge 0,\frac{C-(c^{\star}\wedge 0)}{\varepsilon}]$. Note that $K^{\prime}$ does not depend on $\nu$. We may now let $K=K^{\prime}$ to see that if $\mathcal{A}$ is not tight then $\exists\nu\in\mathcal{A}$ for which $J_{L_{n}}(\nu)>C$. As $\varepsilon$ was arbitrary, the claim follows. ∎
@@
As the loss is lower bounded, we may let this constant be $c:=\inf_{\theta}L(X_{1}^{n},\,\theta)$. Suppose without loss of generality that this constant is negative. Then we claim that $\forall C>c$ the sublevel sets of the GVI functional $T_{n}(\cdot)$ are weak compact. As the loss and the divergence are lower semi–continuous, it follows that $\nu\mapsto J_{L_{n}}(\nu)$ is lower semi–continuous and so $\nu\mapsto T_{n}(\nu):=J_{L_{n}}(\nu)+D(\nu:\Pi)$ is lower semi–continuous [ambrosio2008, dembo2010]. Hence, by Prokhorov’s theorem we need to show that sublevel sets are tight [billingsley1999]. This is immediate if $\Theta$ is compact, so we now show the result for $\Theta$ that are not.
[…]
$\liminf_{k}T_{n}(\nu_{k})\geq T_{n}(\nu_{\star})$
For the uniqueness part, suppose that there exist two minimisers, $\nu_{1}$ and $\nu_{2}$. Then, as $\mathcal{Q}$ is convex we may pick $a\in(0,1)$ and let $\nu:=a\nu_{1}+(1-a)\nu_{2}$ such that
$=aT_{n}(\nu_{1})+(1-a)T_{n}(\nu_{2})$
where the inequality follows by the convexity of the divergence. But as $\nu_{1}$ and $\nu_{2}$ are both minimisers $T_{n}(\nu_{1})=T_{n}(\nu_{2})$ so $\nu$ would further lower the GVI objective, so these cannot be minimisers. Hence, we conclude that the only way $\nu=a\nu_{1}+(1-a)\nu_{2}$ is if $\nu_{1}=\nu_{2}$ (in the sense of measures), and the minimiser may not be written as a convex combination of elements of $\mathcal{Q}$ so this minimiser is an extreme point and unique. ∎
6.2. Characterising the GVI Posterior through the Empirical Loss Minimiser
Lemma 2.
Let Assumptions 1, 2 and 5 hold. Suppose $\exists C\in\mathbb{R}$ and that $\exists P,Q\in\mathcal{Q}$ such that
then $\forall\Pi\in\mathcal{G}$ it holds that $T_{n}(P)<T_{n}(Q)$.
Proof.
For any $P,Q\in\mathcal{Q}$ satisfying the assumption above, we have that:
hence $T_{n}(P)<T_{n}(Q)$ for all priors $\Pi\in\mathcal{G}$. ∎
Proof of Proposition 5.
By Definition 2, $\exists P_{n}^{\star}\in\mathcal{Q}$ such that
$P_{n}^{\star}\in\operatorname*{arg\,min}_{Q\in\mathcal{Q}}\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)].$
Now let $C=\mathbb{E}_{P_{n}^{\star}}[L(X_{1}^{n},\,\theta)]$. Then for all $Q\in\mathcal{Q}$ that satisfy $\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]>C+M(n\beta)^{-1},$ by Lemma 2 it must hold that $T_{n}(P_{n}^{\star})<T_{n}(Q)$, and hence $Q(d\theta)$ is not a minimiser of the GVI objective. Therefore, the only measures $Q^{\prime}\in\mathcal{Q}$ that are potential minimisers of the GVI objective for any prior $\Pi\in\mathcal{G}$ are those that satisfy: $T_{n}(Q^{\prime})\leq T_{n}(P_{n}^{\star})$.